OpenStudy (x3_drummerchick):

can someone explain to me how my teacher got this answer? its using substitution. Will give medals! give me one sec to post the problem

OpenStudy (x3_drummerchick):

\[x^{1/2}+3x^{-1/2}-54x^{-3/2}=0\]

OpenStudy (x3_drummerchick):

solve for x

OpenStudy (anonymous):

\[\sqrt{x}+\frac{ 3 }{ \sqrt{x} }-\frac{ 54 }{ \left( \sqrt{x} \right)^3 }=0\] \[put~\sqrt{x}=t\] \[t+\frac{ 3 }{ t }-\frac{ 54 }{ t^3 }=0\] \[multiply~ by ~t^3\] \[t^4+3t^2-54=0\] put t^2=y \[y^2+3y-54=0\] \[y^2+9y-6y-54=0\] y=? then t=? x=?

OpenStudy (x3_drummerchick):

my teacher told us to try it like this: @surjithayer

OpenStudy (x3_drummerchick):

i dont understand the way he showed us in the above file

Nnesha (nnesha):

yea i was abt to solve like ^^that (Factor )

OpenStudy (x3_drummerchick):

where did the y come from?

OpenStudy (x3_drummerchick):

our teacher wants us to factor it

OpenStudy (x3_drummerchick):

i just dont get how he pulled out that exponential fraction

Nnesha (nnesha):

but i also like the way surjithayer's solved

OpenStudy (x3_drummerchick):

i need to learn how to factor out the negative exponent :(

Nnesha (nnesha):

he supposed that t^2 = y

Nnesha (nnesha):

so t^4 is same as t^2 times t^2 replace t^2 with y y times y = y^2 :=)

OpenStudy (x3_drummerchick):

i know how to use the substitution method, im just a little confused with how x^-1/2 was fatored out.

Nnesha (nnesha):

x^{-1/2) or x^{-3/2} ?

OpenStudy (x3_drummerchick):

x^-3/2, sorry

Nnesha (nnesha):

what's the lowest exponent of x ? that would be your common factor

Nnesha (nnesha):

wait i'll give you an example 2x^2+3x <--x^1 degree would be common factor

OpenStudy (x3_drummerchick):

okay gotcha

Nnesha (nnesha):

so what is common factor in ur equation ?

OpenStudy (x3_drummerchick):

x^-3/2, ause its the smallest

Nnesha (nnesha):

yes right!\[\huge\rm x^\frac{ 1 }{ 2 }+ 3x^\frac{ -1 }{ 2}+54x^\frac{ -3 }{ 2 }\] take out the common factor when you take out x^{-3/2} from x^{1/2} what will ou have left ?\[x^\frac{ -3 }{ 2}(??????????????)\]

OpenStudy (x3_drummerchick):

i dont know

Nnesha (nnesha):

in other words divide all 3 terms by common factor \[\huge\rm \frac{ x^\frac{ 1 }{ 2 } }{ x^\frac{ -3 }{ 2 } }+\frac{ 3x^\frac{ -1 }{ 2 } }{ x^\frac{ -3 }{ 2 } }+\frac{ 54x^\frac{ -3 }{ 2 } }{ x^\frac{ -3 }{ 2 }}\]

Nnesha (nnesha):

remember when we divide same bases we should `subtract `their exponents (exponent rule) \[ \huge\rm \frac{ x^m }{ x^n }=x^{m-n}\]

OpenStudy (x3_drummerchick):

^ thats ringing a bell

Nnesha (nnesha):

:P

Nnesha (nnesha):

example \[\huge\rm \frac{ x^2 }{ x^1}=x^{2-1}\]

OpenStudy (x3_drummerchick):

thats making such more sense

Nnesha (nnesha):

\[\huge\rm \color{ReD}{ \frac{ x^\frac{ 1 }{ 2 } }{ x^\frac{ -3 }{ 2 } }}+\frac{ 3x^\frac{ -1 }{ 2 } }{ x^\frac{ -3 }{ 2 } }+\frac{ 54x^\frac{ -3 }{ 2 } }{ x^\frac{ -3 }{ 2 }}\] look at first term bases are the same so subtract their exponents \[\huge\rm \color{Red}{x^{\frac{ 1 }{ 2 }-(\frac{ -3 }{ 2 })}}\]

OpenStudy (x3_drummerchick):

i think i get it now, ou have to diivide all by it, that reduces it to the ax^2+bx+c

Nnesha (nnesha):

ye!!!

OpenStudy (x3_drummerchick):

wow youre good, thanks!!!!

Nnesha (nnesha):

\[\huge\rm x^\frac{ -3 }{ 2 }(????+???+??)\] write ur answer in the parentheses

OpenStudy (x3_drummerchick):

(x^2+3x-54)

Nnesha (nnesha):

yep right!

OpenStudy (x3_drummerchick):

ive got it from this point out. shew you have no idea how long ive been trying to decifer this shenninagans!

Nnesha (nnesha):

aww se now you got it! great job!

OpenStudy (x3_drummerchick):

thank you so much